Use your financial calculator to determine the monthly payments for each of the following ​$109500 mortgage loans. Assume no prepayments. a. ​30-year fixed at 8.5 percent b.​ 15-year fixed at 7.5 percent c.​ 20-year fixed at 8 percent
Answer #1

a. N = 30 * 12 = 360

I/Y = 8.5/12 = 0.7083333333

PV = -109,500

FV = 0

CPT PMT

PMT = 841.9602645

b. N = 15 * 12 = 180

I/Y = 7.5/12 = 0.625

PV = -109,500

FV = 0

CPT PMT

PMT = 1,015.078534

c. N = 20 * 12 = 240

I/Y = 8/12 = 0.6666666667

PV = -109,500

FV = 0

CPT PMT

PMT = 915.9018755